# PepCompass: Navigating peptide embedding spaces using Riemannian Geometry

### v4

@@ Abstract
Antimicrobial peptide discovery is challenged by the astronomical size of peptide space and the relative scarcity of active peptides. Generative models provide continuous latent “maps” of peptide space, but conventionally ignore decoder-induced geometry and rely on flat Euclidean metrics, rendering exploration and optimization distorted and inefficient. Prior manifold-based remedies assume fixed intrinsic dimensionality, which critically fails in practice for peptide data. Here, we introduce PepCompass, a geometry-aware framework for peptide exploration and optimization. At its core, we define a Union of $\kappa$-Stable Riemannian Manifolds $\mathbb{M}^{\kappa}$, a family of decoder-induced manifolds that captures local geometry while ensuring computational stability. We propose two local exploration methods: Second-Order Riemannian Brownian Efficient Sampling, which provides a convergent second-order approximation to Riemannian Brownian motion, and Mutation Enumeration in Tangent Space, which reinterprets tangent directions as discrete amino-acid substitutions. Combining these yields Local Enumeration Bayesian Optimization (LE-BO), an efficient algorithm for local activity optimization. Finally, we introduce Potential-minimizing Geodesic Search (PoGS), which interpolates between prototype embeddings along property-enriched geodesics, biasing discovery toward seeds, i.e. peptides with favorable activity. In-vitro validation confirms the effectiveness of PepCompass: PoGS yields four novel seeds, and subsequent optimization with LE-BO discovers 25 highly active peptides with broad-spectrum activity, including against resistant bacterial strains. These results demonstrate that geometry-informed exploration provides a powerful new paradigm for antimicrobial peptide design.
Keywords union of manifolds $\cdot$ Peptide optimization $\cdot$ riemannian geometry $\cdot$ geodesics $\cdot$ tangent space $\cdot$ mutation
@@ 1 Introduction
Efficient exploration of peptide space is notoriously difficult. At the global level, there are more than $3.3\times 10^{32}$ combinatorially possible amino acid sequences of length at most 25. At the local level, each peptide of length 25 has nearly 1000 neighbors within an edit radius of one. Moreover, only a small fraction of amino acid sequences correspond to antimicrobial peptides (AMPs), which have high activity against bacteria [48, 49]. This extreme combinatorial complexity renders AMP discovery by brute-force exploration intractable. To address this challenge, we turn to one of the most prolific inventions of humankind: maps.
Since the dawn of civilization, maps have provided a structured way to support both local and global navigation, driving scientific discovery. In modern machine learning, latent-space generative models such as VAEs, GANs, WAEs, and normalizing flows [7] enable building continuous latent representations—maps—of peptides. Such maps have already facilitated the discovery of promising new AMPs [50, 38, 54, 12]. The standard workflow assumes that once the model is trained, its latent space together with the decoder properly models a set of valid, synthetizable peptides. The latent space is typically chosen to be $\mathbb{R}^{d}$ with a flat Euclidean metric, enabling direct application of existing exploration and optimization algorithms. However, such flat representations suffer from a significant flaw: they ignore the differential geometry induced by the decoder, leading to distortions in distances.
Typical approaches attempting to circumvent this problem assume the manifold hypothesis [5] and use the pullback metric [3]. However, recent work [35, 9, 56] has shown that the manifold hypothesis does not withstand empirical scrutiny for image data, where sets of images are better modeled as unions or CW-complexes of manifolds with varying low dimensionality (we refer to Appendix A for Related Work). We show that peptide spaces suffer from a similar issue and introduce decoder-derived Union of $\kappa$-Stable Riemannian Manifolds $\mathbb{M}^{\kappa}$, which captures both the complex structure of peptide space and its local geometry, with computational stability controlled by a parameter $\kappa$. Intuitively, we cut the globally distorted map into a set of charts that enable efficient and distortion-free exploration and optimization.
@@
For local search on a single manifold from the family $\mathbb{M}^{\kappa}$, we designed two geometry-informed approaches: Second-Order Riemannian Brownian Efficient Sampling (SORBES) and Mutation Enumeration in Tangent Space (MUTANG). SORBES is a provably convergent, second-order approximation of the Riemannian Brownian motion [42, 26], serving as a Riemannian analogue of local Gaussian search. MUTANG addresses the discrete nature of peptide space by reinterpreting the local tangent space not as continuous vectors but as discrete mutations, directly corresponding to amino-acid substitutions. This reinterpretation provides both interpretability and efficiency, enabling enumeration of a given peptide’s neighbours. We further combine SORBES and MUTANG into an iterative Local Enumeration procedure that densely populates the neighbourhood of a given peptide with valid, diverse neighbours. Finally, integrating this enumeration with a Bayesian optimization scheme yields an efficient Local Enumeration Bayesian Optimization procedure (LE-BO; see Figure 1B,C).
@@ 2.1 Background
$X=\operatorname{Dec}_{\theta}(\mathcal{Z})\subset\mathbb{R}^{L\times A},\qquad\mathtt{p}(z)=\operatorname{argmax}(\operatorname{Dec}_{\theta}(z),\operatorname{dim}=1)\in\mathcal{A}^{L},$
@@
A common approach is to equip $\mathcal{Z}$ with the standard Euclidean inner product $\langle\cdot,\cdot\rangle_{\mathbb{R}^{d}}$, and use the associated Euclidean distance as a base for exploration. However, this ignores the geometry induced by the decoder, which, under the manifold hypothesis, can be accounted for using the pullback metric [5, 3].
@@
Note that $\hat{\operatorname{Dec}}:(\mathcal{Z},G_{\operatorname{Dec}})\to(\hat{X},\langle\cdot,\cdot\rangle_{\mathbb{R}^{LA}})$ is an isometric diffeomorphism.
@@
Previous work demonstrates that the manifold hypothesis often fails for complex data such as images [35, 9, 56], and the data is better represented as unions of local manifolds of varying dimension, typically lower than that of the latent space. However, the previous methods defined the submanifolds based on pre-specified datasets and could not generalize to new data.
@@ 2.2 Union of $\kappa$-Stable Riemannian Manifolds
Assuming that the manifold hypothesis is indeed violated and that a given generative model has learned to faithfully capture the lower-dimensional structure in the data, it should be reflected in the decoder having rank strictly smaller than the latent dimensionality. We verified this phenomenon for antimicrobial peptides data in two state-of-the-art latent generative models [13, 50] (see Appendix B). This implies that the $G_{Dec}$ is not of full rank, and consequently the pair $(\mathcal{Z},G_{\operatorname{Dec}})$ does not constitute a Riemannian manifold.
@@ 2.3 SORBES - Second-Order Riemannian Brownian efficient Sampling
1: Initialize $z^{\epsilon}_{0}\leftarrow z$, stopped $\leftarrow$ False, $\sigma\leftarrow 0$ ($\sigma$ tracks diffusion time)
@@
4:  Sample a unit tangent direction $\overline{v}\in S_{z}^{\kappa}=\{u\in T_{z}M_{z}^{\kappa}:\langle u,u\rangle^{\operatorname{Dec}}_{z}=1\}$
@@
7:   Update ($\Gamma$ denotes the local Christoffel symbol for $M_{z}^{\kappa}$)
@@
15:end for
16:return $(z^{\epsilon}_{i})_{0\leq i\leq\lfloor\frac{T}{\epsilon^{2}}\rfloor},\ \sigma$
@@
For the proof, see Appendix D. Intuitively, in the small–step limit, our algorithm converges to Riemannian Brownian motion on $M_{z}^{\kappa}(\alpha)$, stopped at the boundary, with the deviation probability decaying exponentially in the inverse time horizon. Theorem 1 extends the main convergence result of Schwarz et al. [42] to possibly non-compact manifolds. Importantly, Schwarz et al. [42] showed that achieving this convergence requires a second-order correction term (capturing the effect of Christoffel symbols), rather than the commonly used naive first-order update. This motivates our use of the second-order scheme in Algorithm 1. In Appendix E (Algorithm 4), we describe SORBES-SE (Stable/Efficient), an implementation of SORBES that approximates the second-order correction using finite differences. It employs an adaptive step size $\epsilon$ that adjusts to the local curvature of the space, while still preserving the convergence guarantees.
@@ 2.4 MUTANG - Mutation Enumeration in Tangent Space
The details of MUTANG are provided in Appendix F (Algorithm 5).
Figure 2: Tangent space as mutation space and local enumeration. A) Around an example peptide GTP we consider two orthogonal peptide-space tangent directions $\Delta\operatorname{Dec}^{(j)}$ obtained from the SVD of the decoder Jacobian at the peptide code. Each direction suggests a specific substitution: $\texttt{T}_{1}\!\to\!\texttt{K}$ and $\texttt{P}_{2}\!\to\!\texttt{C}$. B) Each $\Delta\operatorname{Dec}^{(j)}$ is reshaped into an $L\times A$ map (rows: amino acids; columns: positions). C) Thresholded entries define per-position sets of admissible residues (identity always included). D) The candidate set is the Cartesian product $\mathcal{C}(\texttt{GTP})=\prod_{\ell}S_{\ell}$.
@@ 2.5 Local enumeration
We aim to densely populate the neighbourhood of a single prototype peptide with valid, diverse candidates. Starting from a seed peptide $\mathtt{p}$ with latent code $z_{0}$, we launch multiple Riemannian random walk trajectories using the SORBES algorithm (Sec. 2.3). At the start and after each step, the current latent state is decoded into a peptide and augmented with additional variants generated by MUTANG (Sec. 2.4). Additionally, to account for local dimension variability, we re-estimate the $\kappa$-stable submanifold $M^{\kappa}_{z}$ at each step of the random walk. The union of all decoded walk steps and tangent-space mutations yields a compact, high-quality local candidate set around $\mathtt{p}$. LocalEnumeration is presented in Algorithm 2.
@@
0: seed peptide $\mathtt{p}$, $\kappa_{\textsc{SORBES}},\kappa_{\textsc{MUTANG}}\geq 0$, number of trajectories $M$, walk time budget $T_{\mathrm{walk}}$, nominal step size $\epsilon$, mutation threshold $\theta_{\text{mut}}$
@@
3:  $z\leftarrow z_{0}$; $t\leftarrow 0$
4:  $\mathcal{C}\leftarrow\mathcal{C}\ \cup\ \textsc{MUTANG}(z,\kappa_{\textsc{MUTANG}},\theta_{\text{mut}})$
5:  while $t<T_{\mathrm{walk}}$ do
6:   $((\_,z),\sigma)\leftarrow\textsc{SORBES-SE}(z,\kappa_{\textsc{SORBES}},\epsilon,\texttt{STEP}_{\texttt{max}}{=}1)$ (single step of SORBES-SE)
7:   $t\leftarrow t+\sigma^{2}$; $\mathcal{C}\leftarrow\mathcal{C}\cup\{\mathtt{p}(z)\}$
8:   $\mathcal{C}\leftarrow\mathcal{C}\ \cup\ \textsc{MUTANG}(z,\kappa_{\textsc{MUTANG}},\theta_{\text{mut}})$
9:  end while
10:end for
11:return $\mathcal{C}$ (local candidate set)
@@ 2.6 LE-BO - Local Enumeration Bayesian Optimization
Finally, we integrate our LocalEnumeration algorithm into a Bayesian optimization [21] framework for peptide design, which we term Local Enumeration Bayesian Optimization (LE-BO). Instead of performing costly optimization of the acquisition function in the latent space, which typically relies on Euclidean-distance kernels and ignores both the latent geometry and the discrete nature of peptides, we use surrogate Gaussian process models [43], defined directly in the peptide space. The acquisition function is optimized by locally enumerating peptides in the vicinity of the most promising candidates and then selecting the peptide that maximizes the acquisition value. To further encourage exploration and increase the diversity of discovered candidates, we employ the ROBOT scheme [37], which promotes searching across a broader set of promising regions. Details of LE-BO are presented in Algorithm 3.
@@
3:  $\mathcal{D}:=\mathcal{D}\cup\textsc{LocalEnumeration}(\mathtt{p}_{\text{current}})$ Explore the neighborhood of $\mathtt{p}_{\text{current}}$
4:  $\operatorname{GP}:=\operatorname{GP.fit}(\mathcal{E})$ Fit the surrogate GP model
5:  $\mathcal{D}_{\text{trust}}:=\{\mathtt{p}\in\mathcal{D}\mid\text{Levenshtein}(\mathtt{p},\mathtt{p}_{\text{best}})\leq d_{\text{trust}}\}$ Define the trust region
@@
7:   $\mathtt{p}_{\text{ROBOT}}^{i}:=\operatorname*{arg\,max}_{\mathtt{p}\in\mathcal{D}_{\text{trust}}}\operatorname{GP.acquistion}(\mathtt{p})$
@@
9:   $\mathcal{D}_{\text{trust}}:=\mathcal{D}_{\text{trust}}\setminus\{\mathtt{p}\in\mathcal{D}\mid\text{Levenshtein}(\mathtt{p},\mathtt{p}^{i}_{\text{ROBOT}})\leq d_{\text{ROBOT}}\}$ ROBOT diversity filtering
10:  end for
11:  $p_{\text{current}}:=\arg\min_{1\leq i\leq k_{\text{ROBOT}}}\textsc{Oracle}(\mathtt{p}_{\text{ROBOT}}^{i})$
12:  if $\textsc{Oracle}(\mathtt{p}_{\text{current}})\geq\textsc{Oracle}(\mathtt{p}_{\text{best}})$ then
13:   $\mathtt{p}_{\text{best}}:=\mathtt{p}_{\text{current}}$
14:  end if
15:end for
16:return $\mathtt{p}_{\text{best}}$

### v5

@@ Abstract
Antimicrobial peptide discovery is challenged by the astronomical size of peptide space and the relative scarcity of active peptides. While generative models provide latent maps of this space, they typically ignore decoder-induced geometry and rely on flat Euclidean metrics, making exploration distorted and inefficient. Existing manifold-based approaches assume fixed intrinsic dimensionality, which fails for real peptide data. We introduce PepCompass, a geometry-aware framework based on a Union of $\kappa$-Stable Riemannian Manifolds that captures local decoder geometry while maintaining computational stability. PepCompass performs global interpolation via Potential-minimizing Geodesic Search (PoGS) to bias discovery toward promising seeds, and enables local exploration through Second-Order Riemannian Brownian Efficient Sampling and Mutation Enumeration in Tangent Space, which together form Local Enumeration Bayesian Optimization (LE-BO). PepCompass achieves a 100% in-vitro validation rate: PoGS identifies four novel seeds and LE-BO optimizes them into 25 highly active, broad-spectrum peptides, demonstrating that geometry-informed exploration is a powerful paradigm for antimicrobial peptide design.
Machine Learning, ICML
[…]
@@ 1 Introduction
Efficient exploration of peptide space is notoriously difficult. At the global level, there are more than $3.3\times 10^{32}$ combinatorially possible amino acid sequences of length at most 25. At the local level, each peptide of length 25 has nearly 1000 neighbors within an edit radius of one. Moreover, only a small fraction of amino acid sequences correspond to antimicrobial peptides (AMPs), which have high activity against bacteria (Szymczak and Szczurek, 2023; Szymczak et al., 2025). This extreme combinatorial complexity renders AMP discovery by brute-force exploration intractable. To address this challenge, we turn to one of the most prolific inventions of humankind: maps.
Since the dawn of civilization, maps have provided a structured way to support both local and global navigation, driving scientific discovery. In modern machine learning, latent-space generative models such as VAEs, GANs, WAEs, and normalizing flows (Bond-Taylor et al., 2022) enable building continuous latent representations—maps—of peptides. Such maps have already facilitated the discovery of promising new AMPs (Szymczak et al., 2023; Oort et al., 2021; Wang et al., 2023; Das et al., 2021). The standard workflow assumes that once the model is trained, its latent space together with the decoder properly models a set of valid, synthetizable peptides. The latent space is typically chosen to be $\mathbb{R}^{d}$ with a flat Euclidean metric, enabling direct application of existing exploration and optimization algorithms. However, such flat representations suffer from a significant flaw: they ignore the differential geometry induced by the decoder, leading to distortions in distances.
Typical approaches attempting to circumvent this problem assume the manifold hypothesis (Bengio et al., 2014) and use the pullback metric (Arvanitidis et al., 2021). However, recent work (Brown et al., 2023; Loaiza-Ganem et al., 2024; Wang and Wang, 2024) has shown that the manifold hypothesis does not withstand empirical scrutiny for image data, where sets of images are better modeled as unions or CW-complexes of manifolds with varying low dimensionality (we refer to Appendix A for Related Work). We show that peptide spaces suffer from a similar issue and introduce decoder-derived Union of $\kappa$-Stable Riemannian Manifolds $\mathbb{M}^{\kappa}$, which captures both the complex structure of peptide space and its local geometry, with computational stability controlled by a parameter $\kappa$. Intuitively, we cut the globally distorted map into a set of charts that enable efficient and distortion-free exploration and optimization.
@@
For local search on a single manifold from the family $\mathbb{M}^{\kappa}$, we designed two geometry-informed approaches: Second-Order Riemannian Brownian Efficient Sampling (SORBES) and Mutation Enumeration in Tangent Space (MUTANG). SORBES is a provably convergent, second-order approximation of the Riemannian Brownian motion (Schwarz et al., 2023; Herrmann et al., 2023), serving as a Riemannian analogue of local Gaussian search. MUTANG addresses the discrete nature of peptide space by reinterpreting the local tangent space not as continuous vectors but as discrete mutations, directly corresponding to amino-acid substitutions. This reinterpretation provides both interpretability and efficiency, enabling enumeration of a given peptide’s neighbours. We further combine SORBES and MUTANG into an iterative Local Enumeration procedure that densely populates the neighbourhood of a given peptide with valid, diverse neighbours. Finally, integrating this enumeration with a Bayesian optimization scheme yields an efficient Local Enumeration Bayesian Optimization procedure (LE-BO; see Figure 1B,C).
@@ 2.1 Background
$X=\operatorname{Dec}_{\theta}(\mathcal{Z})\subset\mathbb{R}^{L\times A},$
$\mathtt{p}(z)=\operatorname{argmax}(\operatorname{Dec}_{\theta}(z),\operatorname{dim}=1)\in\mathcal{A}^{L},$
@@
A common approach is to equip $\mathcal{Z}$ with the standard Euclidean inner product $\langle\cdot,\cdot\rangle_{\mathbb{R}^{d}}$, and use the associated Euclidean distance as a base for exploration. However, this ignores the geometry induced by the decoder, which, under the manifold hypothesis, can be accounted for using the pullback metric (Bengio et al., 2014; Arvanitidis et al., 2021).
@@
Note that $\hat{\operatorname{Dec}}:(\mathcal{Z},G_{\operatorname{Dec}})\to(\hat{X},\langle\cdot,\cdot\rangle_{\mathbb{R}^{LA}})$ is an isometric diffeomorphism, meaning that distances measured in the latent space locally match distances in ambient space.
@@
Previous work demonstrates that the manifold hypothesis often fails for complex data such as images (Brown et al., 2023; Loaiza-Ganem et al., 2024; Wang and Wang, 2024), and the data is better represented as unions of local manifolds of varying dimension, typically lower than that of the latent space. However, the previous methods defined the submanifolds based on pre-specified datasets and could not generalize to new data.
@@ 2.2 Union of $\kappa$-Stable Riemannian Manifolds
Assuming that the manifold hypothesis is indeed violated and that a given generative model has learned to faithfully capture the lower-dimensional structure in the data, it should be reflected in the decoder having rank strictly smaller than the latent dimensionality. We verified this phenomenon for antimicrobial peptides data in two state-of-the-art latent generative models (Das et al., 2018; Szymczak et al., 2023) (see Appendix B). This implies that the $G_{Dec}$ is not of full rank, and consequently the pair $(\mathcal{Z},G_{\operatorname{Dec}})$ does not constitute a Riemannian manifold.
@@
On the usage of Euclidean metric in the ambient space
Note that in the construction of the $\kappa$-stable manifolds, we endow the decoder output space, that is position-factorized peptide probabilities with the standard Euclidean metric. Although this choice may appear simplistic, it induces a mutationally uniform geometry that naturally penalizes non-local insertions and deletions, which are more likely to induce substantial changes in peptide structure. In Appendix D we provide a formal derivation of these properties and discuss potential extensions to biologically informed ambient metrics.
@@ 2.3 SORBES - Second-Order Riemannian Brownian efficient Sampling
1: Initialize $z^{\epsilon}_{0}\leftarrow z$, stopped $\leftarrow$ False, $\sigma\leftarrow 0$
($\sigma$ tracks diffusion time)
@@
4:  $\overline{v}\in S_{z}^{\kappa}=\{u\in T_{z}M_{z}^{\kappa}:\langle u,u\rangle^{\operatorname{Dec}}_{z}=1\}$
(Sample a unit tangent direction)
@@
7:
@@
(Update - $\Gamma$ denotes the local Christoffel symbol for $M_{z}^{\kappa}$)
@@
15:end forreturn $(z^{\epsilon}_{i})_{0\leq i\leq\lfloor\frac{T}{\epsilon^{2}}\rfloor},\ \sigma$
@@
For the proof, see Appendix E. Intuitively, in the small–step limit, our algorithm converges to Riemannian Brownian motion on $M_{z}^{\kappa}(\alpha)$, stopped at the boundary, with the deviation probability decaying exponentially in the inverse time horizon. Theorem 2.1 extends the main convergence result of Schwarz et al. (2023) to possibly non-compact manifolds. Importantly, Schwarz et al. (2023) showed that achieving this convergence requires a second-order correction term (capturing the effect of Christoffel symbols), rather than the commonly used naive first-order update. This motivates our use of the second-order scheme in Algorithm 1. In Appendix F (Algorithm 4), we describe SORBES-SE (Stable/Efficient), an implementation of SORBES that approximates the second-order correction using finite differences. It employs an adaptive step size $\epsilon$ that adjusts to the local curvature of the space, while still preserving the convergence guarantees.
@@ 2.4 MUTANG - Mutation Enumeration in Tangent Space
The details of MUTANG are provided in Appendix G (Algorithm 5).
Figure 2: Tangent space as mutation space and local enumeration. A) For an example peptide GTP we consider two orthogonal peptide-space tangent directions $\Delta\operatorname{Dec}^{(j)}$ obtained from the SVD of the decoder Jacobian. Each direction suggests a specific substitution: $\texttt{T}_{1}\!\to\!\texttt{K}$ and $\texttt{P}_{2}\!\to\!\texttt{C}$. B) Each $\Delta\operatorname{Dec}^{(j)}$ is reshaped into an $L\times A$ map (rows: amino acids; columns: positions). C) Thresholded entries define per-position sets of admissible residues (identity always included). D) The candidate set is the Cartesian product $\mathcal{C}(\texttt{GTP})=\prod_{\ell}S_{\ell}$.
@@ 2.5 Local enumeration
We aim to densely populate the neighbourhood of a single prototype peptide with valid, diverse candidates. Starting from a seed peptide $\mathtt{p}$ with latent code $z_{0}$, we launch multiple Riemannian random walk trajectories using the SORBES algorithm (Section 2.3). At the start and after each step, the current latent state is decoded into a peptide and augmented with additional variants generated by MUTANG (Section 2.4). Additionally, to account for local dimension variability, we re-estimate the $\kappa$-stable submanifold $M^{\kappa}_{z}$ at each step of the random walk. The union of all decoded walk steps and tangent-space mutations yields a compact, high-quality local candidate set around $\mathtt{p}$. Despite leveraging second-order geometric information, LocalEnumeration remains computationally efficient: execution time of each step scales linearly with the latent dimension, peptide length, and decoder evaluation time. Algorithm 2 summarizes the procedure, with a detailed time and memory complexity analysis provided in Appendix H.
@@
0: seed peptide $\mathtt{p}$, $\kappa_{\textsc{SORBES}},\kappa_{\textsc{MUTANG}}\geq 0$, number of trajectories $M$, walk time $T_{\mathrm{walk}}$, step size $\epsilon$, mutation threshold $\theta_{\text{mut}}$
@@
3:  $z\leftarrow z_{0}$; $t\leftarrow 0$; $\mathcal{C}\leftarrow\mathcal{C}\ \cup\ \textsc{MUTANG}(z,\kappa_{\textsc{MUTANG}},\theta_{\text{mut}})$
4:  while $t<T_{\mathrm{walk}}$ do
5:   $((\_,z),\sigma)\leftarrow\textsc{SORBES-SE}(z,\kappa_{\textsc{SORBES}},\epsilon,\texttt{STEP}_{\texttt{max}}{=}1)$ (single step of SORBES-SE)
6:   $t\leftarrow t+\sigma^{2}$; $\mathcal{C}\leftarrow\mathcal{C}\cup\{\mathtt{p}(z)\}$
7:   $\mathcal{C}\leftarrow\mathcal{C}\ \cup\ \textsc{MUTANG}(z,\kappa_{\textsc{MUTANG}},\theta_{\text{mut}})$
8:  end while
9:end forreturn $\mathcal{C}$ (local candidate set)
@@ 2.6 LE-BO - Local Enumeration Bayesian Optimization
Finally, we integrate our LocalEnumeration algorithm into a Bayesian optimization (Garnett, 2023) framework for peptide design, which we term Local Enumeration Bayesian Optimization (LE-BO). Instead of performing costly optimization of the acquisition function in the latent space, which typically relies on Euclidean-distance kernels and ignores both the latent geometry and the discrete nature of peptides, we use surrogate Gaussian process models (Seeger, 2004), defined directly in the peptide space. The acquisition function is optimized by locally enumerating peptides in the vicinity of the most promising candidates and then selecting the peptide that maximizes the acquisition value. To further encourage exploration and increase the diversity of discovered candidates, we employ the ROBOT scheme (Maus et al., 2023), which promotes searching across a broader set of promising regions. Details of LE-BO are presented in Algorithm 3.
@@
3:  $\mathcal{D}:=\mathcal{D}\cup\textsc{LocalEnumeration}(\mathtt{p}_{\text{current}})$
(Explore the neighborhood of $\mathtt{p}_{\text{current}}$)
4:  $\operatorname{GP}:=\operatorname{GP.fit}(\mathcal{E})$ (Fit the surrogate GP model)
5:  $\mathcal{D}_{\text{trust}}:=\{\mathtt{p}\in\mathcal{D}\mid\text{Levenshtein}(\mathtt{p},\mathtt{p}_{\text{best}})\leq d_{\text{trust}}\}$ (Define the trust region)
@@
7:   $\mathtt{p}_{\text{ROBOT}}^{i}:=\operatorname{argmax}_{\mathtt{p}\in\mathcal{D}_{\text{trust}}}\operatorname{GP.acquistion}(\mathtt{p})$
@@
9:   $\mathcal{D}_{\text{ROBOT}}=\{\mathtt{p}\in\mathcal{D}\mid\text{Levenshtein}(\mathtt{p},\mathtt{p}^{i}_{\text{ROBOT}})\leq d_{\text{ROBOT}}\}$
10:   $\mathcal{D}_{\text{trust}}:=\mathcal{D}_{\text{trust}}\setminus\mathcal{D}_{\text{ROBOT}}$
(ROBOT diversity filtering)
11:  end for
12:  $p_{\text{current}}:=\arg\min_{1\leq i\leq k_{\text{ROBOT}}}\textsc{Oracle}(\mathtt{p}_{\text{ROBOT}}^{i})$
13:  if $\textsc{Oracle}(\mathtt{p}_{\text{current}})\geq\textsc{Oracle}(\mathtt{p}_{\text{best}})$ then
14:   $\mathtt{p}_{\text{best}}:=\mathtt{p}_{\text{current}}$
15:  end if
16:end for
17:return $\mathtt{p}_{\text{best}}$